OpenStudy (anonymous):

solve using substitution: 4x+6y=-9 -3x-10y=-11

OpenStudy (saifoo.khan):

Hey, someoneelse-man!

OpenStudy (anonymous):

waz up seafood!!!

OpenStudy (saifoo.khan):

do u want the answers?

OpenStudy (anonymous):

kk i need help..i hate tese problems, can u show me how to do them

OpenStudy (saifoo.khan):

Dont sleep in class.

OpenStudy (anonymous):

i do not sleep in class!! ma dumb teacher doesnt give us notes, jus tells us to look it up in our text books

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=uzyd_mIJaoc

OpenStudy (anonymous):

this is harder than that.. i dunno wat to plug in...

hero (hero):

4x+6y=-9 -3x-10y=-11 y = (-9 - 4x)/6 y = (-11 + 3x)/-10 y = y -10(-9-4x) = 6(-11+ 3x) 90 + 40x = -66+18x 40x-18x = -66-90 22x = -156 x = -156/22 = -78/11 Good luck finding y

OpenStudy (saifoo.khan):

solve the first equation for x: 4x+6y=-9 x = (-9 -6y)/y Now, substitute this equation in eq2: -3x-10y=-11 -3( (-9 -6y)/y)-10y=-11